OpenStudy (anonymous):

Find three consecutive numbers whose sum is 252.

OpenStudy (anonymous):

83,84,85

OpenStudy (anonymous):

Lets see.. three consecutive nmbers can be thought of as x, x+1, and x+2

OpenStudy (anonymous):

oh k

OpenStudy (anonymous):

Take the summation and set that equal to the number you want to solve for. 252. Then solve for x

OpenStudy (anonymous):

like this x+1=252

OpenStudy (anonymous):

Let x be the first number x + (x+1) + (x+2) = 252 x+x+1+x+2=252 3x+3=252 3x=252-3 3x=249 x=83 first number = 83 next number = x+1 = 84 third number = x+2 = 85

OpenStudy (anonymous):

thank. help with this one Find three consecutive even integers whose sum is 72.

OpenStudy (anonymous):

Use the same mehod

OpenStudy (anonymous):

let x be first even integer x+ (x+2) + (x+4) and the same method follows

OpenStudy (anonymous):

try to work it out realizing "x" will be the value in question.

OpenStudy (anonymous):

so the first number is gonna be 22